OpenStudy (anonymous):

A police officer was sent to investigate a car accident. He knows the speed of the car, s, can be measured with the equation s = at, where a is the rate of acceleration and t is the time. Which equation can he use to solve for the rate of acceleration? a = s – t a = t – s a = s ÷ t a = st My answer is a

OpenStudy (anonymous):

If \[s = at\] and we divide each side by "t", then \[s/t = at/t\] the "t" on the right will cancel out and the final formula will become \[s/t = a\]